OpenStudy (anonymous):

The graphs of four exponential functions are shown below: (attacked below) Which pair of graphs should be plotted to show the solution to (equation below w/ graph)

OpenStudy (anonymous):

\[4(2)^x = 4\left(\begin{matrix}1 \\ 2\end{matrix}\right)^x\]

OpenStudy (aum):

Put x = 0, What is the y-value of \(4(2)^x\) and \(4\left(\frac 12\right)^x\) ?

OpenStudy (anonymous):

4 = 4

OpenStudy (anonymous):

Does that mean A to B? @aum

OpenStudy (aum):

Yes. A and B. B is the increasing function and it represents \(4(2)^x\). A is a decreasing function and represents \(4\left(\frac 12\right)^x\). Both have y-intercepts of 4 which makes it easy to identify in the graph.

OpenStudy (anonymous):

Oh! Thank you! That was stupid easy... can't believe I had to ask.

OpenStudy (aum):

You are welcome.